Consider the following.

t

0

1

2

3

L(t)

775

558

401.76

289.2672

Is the function exponential?

Yes

No    

If the function is exponential, find its equation. (If an answer does not exist, enter DNE.)

L(t) =

Use the equation to predict the next entry to extend its table of values. (Round your answer to six decimal places. If an answer does not exist, enter DNE.)

t

0

1

2

3

4

L(t)

775

558

401.76

289.2672
Answer #1

The given function is exponential.

since, 558/775 = 401.76/558 = 289.2672/401.76 = 0.72

since the ratio is constant we can it is exponential distribution.

to find next one,

x/289.2672 = 0.72

x= 0.72*289.2672= 208.272384.

therfore L(4)= 208.272384